Find the mass of water that vaporizes when 4.16 kg of mercury at 224 °C is added to 0.447 kg of water at 89.8 °C. Help pleaseFind the mass of water that vaporizes when 4.16 kg of mercury at 224 oC is added to 0.447 kg of water at 89.8 °C. UnitšT kg Number TT0.0202 the tolerance is +/-296
Answer #1

Amount of heat lost by mercury,

Q = mc\DeltaT

Q = 4.16 * 140 * (224 - 100)

Q = 72217.6 J

Heat required to change the temperature of water from 89.9 to 100 oC,

Q1 = mc\DeltaT = 0.447*4200*(100 - 89.8)

Q1 = 19149.48 J

Heat needed for phase change = Q2 = mL

Here, Q = Q1 + Q2

72217.6 = 19149.48 + m*2.26*10^6

m = 0.0234 kg